OpenStudy (anonymous):

solve by elimination method 0.3x-0.2y=4 and 0.2x+0.5y=-21/23

OpenStudy (anonymous):

Is -21/23 suppose to be a fraction?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lets get rid of the x first, so we need to multiply first equation by 5, and second by 2 1.5x-.1=20 .4x+.1y=-42/23 Answer for that= 1.9x= 418/23 x is about 9.56522 then plug the x we found into either of the equation and solve for y.